OpenStudy (anonymous):

93.618 divided by .03

OpenStudy (anonymous):

Calculator?

OpenStudy (anonymous):

93.618 /0.03 = 3120.6

OpenStudy (anonymous):

@above Your expression calculates to a value of 2100 according to wolf. http://m.wolframalpha.com/input/?i=9%2F.03+%2B+36%2F.03+%2B+18%2F.03&x=0&y=0 .03 * 2100 = 63, not 93.618

OpenStudy (anonymous):

312.06

OpenStudy (anonymous):

\[\frac{93.618}{.03}=\frac{93618}{30}=3120\frac{18}{30} \]

OpenStudy (anonymous):

I would look at this calculation in a different way. If we make the 0.03 into a whole number by multiplying by 100 = 3 Take the large number and do the same 93.618 x 100 = 9361.8 Now the calculation is easy to do 9361.8 / 3 = 3120.6